OpenStudy (anonymous):

For the function f(x) = 3(x − 1)2 + 2, identify the vertex, domain, and range. The vertex is (1, 2), the domain is all real numbers, and the range is y ≥ 2. The vertex is (1, 2), the domain is all real numbers, and the range is y ≤ 2. The vertex is (–1, 2), the domain is all real numbers, and the range is y ≥ 2. The vertex is (–1, 2), the domain is all real numbers, and the range is y ≤ 2.

OpenStudy (anonymous):

correct answer gets medal and fan

OpenStudy (anonymous):

@Robert136

OpenStudy (anonymous):

@ali2x2

OpenStudy (anonymous):

@heretohelpalways

OpenStudy (anonymous):

@taramgrant0543664

jigglypuff314 (jigglypuff314):

Hi there :) Sorry for being a little late to respond .-. they have given you the equation in "vertex form" where \( f(x) = a(x-h)^2+k \) where the point \( (h,~k) \) is the vertex comparing this "form" to your \( f(x) = 3(x-1)^2 + 2 \) we can see that \( h=1 \) and \(k=2\) putting this as a point \((h,k)\), you would get the proper vertex point that they are asking for :)

jigglypuff314 (jigglypuff314):

The other deciding factor in your options, it seems, would be the range. If you have seen a picture of a \(parabola\) before [it looks like a U] from this we can say that anything above the vertex (in this case) would be included in the range so since we got that the y value of the vertex to be \(k=2\) we can say that the range of this function would be greater than and including \(2\)